# Thermodynamic nonequilibrium effects in three-dimensional high-speed compressible flows: Multiscale modeling and simulation via the discrete Boltzmann method

### v1

@@ Abstract
Three-dimensional (3D) high-speed compressible flow is a typical nonlinear, nonequilibrium, and multiscale complex flow. Traditional fluid mechanics models, based on the quasi-continuum assumption and near-equilibrium approximation, are insufficient to capture significant discrete effects and thermodynamic nonequilibrium effects (TNEs) as the Knudsen number increases. To overcome these limitations, a discrete Boltzmann modeling and simulation method, rooted in kinetic and mean-field theories, has been developed. By applying Chapman-Enskog multiscale analysis, the essential kinetic moment relations $\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{20},\mathbf{M}_{2},\mathbf{M}_{31},\mathbf{M}_{3},\mathbf{M}_{42},\mathbf{M}_{4},\mathbf{M}_{53})$ for characterizing second-order TNEs are determined. These relations $\bm{\Phi}$ are invariants in coarse-grained physical modeling, providing a unique mesoscopic perspective for analyzing TNE behaviors. A discrete Boltzmann model, accurate to the second-order in the Knudsen number, is developed to enable multiscale simulations of 3D supersonic flows. As key TNE measures, nonlinear constitutive relations (NCRs), are theoretically derived for the 3D case, offering a constitutive foundation for improving macroscopic fluid modeling. The NCRs in three dimensions exhibit greater complexity than their two-dimensional counterparts. This complexity arises from increased degrees of freedom, which introduce additional kinds of nonequilibrium driving forces, stronger coupling between these forces, and a significant increase in nonequilibrium components. At the macroscopic level, the model is validated through several classical test cases, ranging from 1D to 3D scenarios, from subsonic to supersonic regimes. At the mesoscopic level, the model accurately captures typical TNEs, such as viscous stress and heat flux, around mesoscale structures, across various scales and orders. This work provides kinetic insights that advance multiscale simulation techniques for 3D high-speed compressible flows.
@@ I Introduction
Beyond applications in aerospace, energy technology, and weapon systems, exploring high-speed compressible flow mechanisms helps overcome the limitations of traditional hydrodynamic models in small-scale, high-pressure conditions. For instance, in high-performance chip cooling within micro-electromechanical systems (MEMS), high-speed gas flowing through micro-nozzles generates localized jets on the chip surface. These jets enhance convective heat transfer by disrupting the thermal boundary layer through shear forces. Additionally, compressible effects at the nozzle exit generate expansion waves, reducing gas temperature, improving cooling efficiency, and mitigating shock wave effectsHo and Tai (1998); Stone, Stroock, and Ajdari (2004). In heap leaching with microporous seepage, high-pressure fluids moving through micropores and fractures induce local pressure and velocity fluctuations. These fluctuations enhance pore wall dissolution and facilitate fracture propagation. Shock waves at pore tips further assist in forming interconnected fracture pathwaysWu, Yao, and Wang (2012). These examples highlight the critical role of high-speed compressible flow in small-scale nonequilibrium processes, where the interplay of high pressures, localized effects, and compressibility drives significant physical and engineering effects.
@@
Due to the complexity of the Boltzmann equation’s collision term, developing simplified kinetic models that capture its most relevant features is an effective approach. Common simplified models include the Bhatnagar-Gross-Krook (BGK) modelBhatnagar, Gross, and Krook (1954), the ellipsoidal statistical BGK (ES-BGK) modelHolway (1966), the Shakhov modelShakhov (1968), and the Liu modelLiu (1990). To further enhance computational efficiency and applicability, researchers have explored various strategies for simplifying collision terms and tailoring models to specific problems. This has resulted in the development of various computational methods for solving cross-scale flow problems. Notable examples include the gas-kinetic schemeXu and Prendergast (1994); Xu (2001), unified gas-kinetic schemeXu and Huang (2010); Huang, Xu, and Yu (2012), discrete unified gas-kinetic schemeGuo, Xu, and Wang (2013); Xin, Zhang, and Guo (2023), discrete velocity methodYang et al. (2016); Yuan and Zhong (2020); Yang et al. (2022), and the lattice Boltzmann method (LBM)He and Luo (1997); Gan et al. (2008, 2011a); Xu et al. (2012); Wang, Wei, and Qian (2022); Shirsat, Nayak, and Patil (2022); Fei et al. (2023); Liang et al. (2024); Wang et al. (2017); Shan et al. (2024); Zhuang et al. (2024); Tao, Guo, and Wang (2017); Chen et al. (2024a).
@@
Historically, the DBM evolved from the physical modeling branch of the LBM through selective enhancements and simplifications. As a discretized form of the simplified Boltzmann equation, DBM retains the use of discrete velocities but avoids relying on specific discrete formats, imposing only the essential physical constraints during discretization. These constraints include conserved kinetic moments and higher-order nonconserved moments, which are invariants in coarse-grained physical modeling and must be preserved. Such constraints provide a unique perspective for analyzing TNE behaviors, effects, and their manifestations. As the order and variety of higher-order kinetic moment constraints increase, DBM goes beyond the continuity assumption and near-equilibrium approximations of traditional fluid modeling. Although the model’s complexity increases slightly, its ability to describe nonequilibrium phenomena is significantly enhanced. Over time, DBM has developed new schemes for detecting, presenting, describing, and analyzing nonequilibrium states and effects based on the phase space description method and complex physical field analysis techniquesXu, Zhang, and Gan (2024). Specifically, each nonconserved kinetic moment provides a perspective on TNE intensity. These perspectives are interrelated and complementary, allowing for a more refined characterization of the physical structure and evolution of mesoscopic interfaces.
@@
While these studies advanced the exploration of TNEs in 3D cases, they are limited to situations with weak nonequilibrium effects. In scenarios involving large gradients or intense TNEs, these models often fail to capture the finer details of the nonequilibrium phenomena. To address this limitation, extending DBM to the 3D Burnett level is essential for achieving a more comprehensive characterization of nonequilibrium properties. In the field of higher-order DBM, Gan et al. derived expressions for nonlinear viscous stress and heat flux by considering second-order TNEs using CE analysis. They constructed a Burnett-level DBM to investigate HNEs and TNEs in high-speed compressible flowsGan et al. (2018a). Later, they developed a multiscale, multiphase DBM based on density functional kinetic theory. This model, corresponding to the super-Burnett level, is suitable for studying TNEs in systems ranging from continuum to transition flow regimesGan et al. (2022).
Building upon previous work, this study constructs a 3D Burnett-level DBM to investigate HNE and TNEs in high-speed compressible flows. The core concept involves applying CE multiscale analysis to identify the kinetic moment relationships that should be preserved in coarse-grained physical modeling when considering second-order TNEs. Additionally, this study provides a theoretical derivation of nonlinear constitutive relations for the 3D case, analyzing how spatial dimensions influence the types of nonequilibrium driving forces and their coupling effects. To facilitate this, a suitable discrete velocity set is designed to discretize the phase space, enabling the development of a robust and efficient three-dimensional (3D) DBM. The structure of the paper is organized as follows: Section II outlines the construction of the multiscale model. Section III validates the model’s accuracy in capturing large-scale structures in high-speed compressible flows, using standard test cases from 1D to 3D scenarios. Section IV evaluates the model’s ability to describe higher-order TNEs across scales. Finally, Section V concludes with a summary of the key findings.
@@ II Three-dimensional high-order discrete Boltzmann model
(2) Analysis stage: By utilizing higher-order nonconserved moments $(f-f{{}^{eq}})$, DBM enables the efficient extraction of nonequilibrium states, effects, and behaviors. Moreover, constructing the phase space through the independent component of $\mathbf{M}_{mn}(f-f{{}^{eq}})$ offers a more intuitive representation of nonequilibrium states, effects, and behaviors.
@@ II.1 Simplification of the Boltzmann equation
In kinetic theory, the properties of a system are described by the distribution function and its kinetic moments. After discretization, the distribution function itself loses physical significance, while the integral or summation forms of the corresponding kinetic moments retain their physical relevance. Therefore, whether simplifying the collision term or discretizing the particle velocity space, the DBM must ensure that the fundamental physical constraint describing the system’s behavior—represented by the kinetic moments—remains invariant, both before discretization in the integral form and after discretization in the summation form, as expressed below:
@@
The CE multiscale analysis demonstrates that, recovering the hydrodynamic equations at different levels, requires satisfying specific kinetic moments. To determine the required kinetic moments, the CE expansion is applied to both sides of Eq. (4), beginning with a multiscale expansion of the distribution function, time derivatives, and spatial derivatives:
@@
where $f_{i}^{(j)}$ represents the $j$-th order deviation from $f_{i}^{(0)}$, and $\partial_{t_{j}}$ denotes the $j$-th order term in the time scale expansion.
@@
It is evident that both $f_{i}^{(1)}$ and $f_{i}^{(2)}$ can be expressed in terms of $f_{i}^{(0)}$. Specifically, $f_{i}^{(1)}$ is represented by a first-order polynomial in $\mathbf{v}_{i}$, while $f_{i}^{(2)}$ is represented by a second-order polynomial in $\mathbf{v}_{i}$. Consequently, to satisfy the moment relations for $f_{i}^{(0)}$, the required accuracy must increase by one order for $f_{i}^{(1)}$ and by two orders for $f_{i}^{(2)}$. Thus, Eq. (5) can be rewritten as:
@@
$\mathbf{M}_{20}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})=\frac{1}{2}\rho[(n+3)RT+u^{2}],$ (14)
@@
$\mathbf{M}_{31}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}=\frac{1}{2}\rho\mathbf{u}[(n+5)RT+u^{2}],$ (16)
@@
$\mathbf{M}_{42}=$ $\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}\mathbf{v}_{i}=\frac{1}{2}\rho[(n+5)R^{2}T^{2}$ (18)
@@
$\mathbf{M}_{53}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}\mathbf{v}_{i}\mathbf{v}_{i}=\rho\left[\left(\frac{n+9}{2}RT+\frac{u^{2}}{2}\right)\mathbf{uuu}\right.$ (20)
@@
$\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{20},\cdots,\mathbf{M}_{4},\mathbf{M}_{53})^{T}=(M_{0},M_{1x},M_{1y},M_{1z},\cdots,M_{53yzz},M_{53zzz})^{T}$ is a $55\times 1$ column matrix. Each element of $\bm{\Phi}$ represents an independent kinetic moment of $f_{i}^{(0)}$. The matrix $\mathbf{C}=(\mathbf{c}_{1},\mathbf{c}_{2},\cdots,\mathbf{c}_{55})$ is a $55\times 5$ 55 square matrix that establishes the relationship between the discrete equilibrium distribution function and the kinetic moments. Here, $\mathbf{c}_{i}$ is the $i$-th column of $\mathbf{C}$, defined as: $\mathbf{c}_{i}=[1,v_{ix},v_{iy},v_{iz},\cdots,\frac{1}{2}(v_{i}^{2}+\eta_{i}^{2})v_{iz}v_{iz}v_{iz}]^{T}$, and $\mathbf{f}^{(0)}=(f_{1}^{(0)},f_{2}^{(0)},f_{3}^{(0)},\cdots,f_{55}^{(0)})^{T}$ represents the vector of discrete equilibrium distribution functions. Thus, $f_{i}^{(0)}$ can be computed as follows Gan et al. (2013):
@@
To efficiently handle the continuity of velocity space in numerical simulations, the originally continuous velocity space is discretized into a finite set of discrete velocities. These nine kinetic moments lead to 55 component equations, requiring at least 55 discrete velocities for accurate modeling. To ensure computational stability and accuracy, a 3D discrete velocity set with fifty-five velocities (D3V55), as shown in Fig. 1(a), is proposed. The mathematical expression is as follows:
$(v_{ix},v_{iy})=\begin{cases}(0,0,0)&\text{for }i=1,\\ cyc:c(\pm 1,0,0)&\text{for }2\leq i\leq 7,\\ cyc:c(\pm 1,\pm 1,0)&\text{for }8\leq i\leq 19,\\ c(\pm 1,\pm 1,\pm 1)&\text{for }20\leq i\leq 27,\\ cyc:c(\pm 1,\pm 2,0)&\text{for }28\leq i\leq 51,\\ c(3,-2,-1)\text{ or }c(-2,1,2)&\text{for }i=52,\\ c(-3,2,1)\text{ or }c(-2,3,-2)&\text{for }i=53,\\ c(2,1,-3)\text{ or }c(1,-2,-1)&\text{for }i=54,\\ c(-2,-1,3)\text{ or }c(3,-2,1)&\text{for }i=55.\end{cases}$ (23)
Here, "$cyc$" represents cyclic permutation, which generates the corresponding vector by cyclically rearranging the digits and signs in the velocity vector. The decomposition diagram is shown in Fig. 1 (b)-(f). Specifically, (i) the first velocity corresponds to the zero velocity; (ii) the second to seventh velocities correspond to the face-centered velocities; (iii) the eighth to nineteenth velocities correspond to the second diagonal velocities; (iv) the twentieth to twenty-seventh velocities correspond to the cube vertex velocities; (v) the twenty-eighth to fifty-first velocities correspond to the third diagonal velocities; and (vi) the final four velocities are special velocities with insufficient symmetry, used to ensure the invertibility of the velocity matrix. We also set
$\eta=\begin{cases}\eta_{i}=10\eta_{0}&\text{for }i=1,\\ \eta_{i}=\eta_{0}&\text{for }2\leq i\leq 7,\,i=15,\,19,\,26,\\ &\hskip 13.00005pt35\leq i\leq 40,\\ \eta_{i}=\eta_{0}\text{ or }\eta_{i}=0&\text{for }52\leq i\leq 53,\\ \eta_{i}=0&\text{for all other }i.\end{cases}$ (24)

### v2

@@ Abstract
Three-dimensional (3D) high-speed compressible flow is a typical nonlinear, nonequilibrium, and multiscale complex flow. Traditional fluid mechanics models, based on the quasi-continuum assumption and near-equilibrium approximation, are insufficient to capture significant discrete effects and thermodynamic nonequilibrium effects (TNEs) as the Knudsen number increases. To overcome these limitations, a discrete Boltzmann modeling and simulation method, rooted in kinetic and mean-field theories, has been developed. By applying Chapman-Enskog multiscale analysis, the essential kinetic moment relations $\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{2,0},\mathbf{M}_{2},\mathbf{M}_{3,1},\mathbf{M}_{3},\mathbf{M}_{4,2},\mathbf{M}_{4},\mathbf{M}_{5,3})$ for characterizing second-order TNEs are determined. These relations $\bm{\Phi}$ are invariants in coarse-grained physical modeling, providing a unique mesoscopic perspective for analyzing TNE behaviors. A discrete Boltzmann model, accurate to second-order in the Knudsen number, is developed to enable multiscale simulations of 3D supersonic flows. As key TNE measures, nonlinear constitutive relations (NCRs), are theoretically derived for the 3D case, offering a constitutive foundation for improving macroscopic fluid modeling. The NCRs in three dimensions exhibit greater complexity than their two-dimensional counterparts. This complexity arises from increased degrees of freedom, which introduce additional kinds of nonequilibrium driving forces, stronger coupling between these forces, and a significant increase in nonequilibrium components. At the macroscopic level, the model is validated through a series of classical test cases, ranging from one-dimensional to 3D scenarios, from subsonic to supersonic regimes. At the mesoscopic level, the model accurately captures typical TNEs, such as viscous stress and heat flux, around mesoscale structures, across various scales and orders. This work provides kinetic insights that advance multiscale simulation techniques for 3D high-speed compressible flows.
@@ I Introduction
Beyond applications in aerospace, energy technology, and weapon systems, exploring high-speed compressible flow mechanisms helps overcome the limitations of traditional hydrodynamic models in small-scale, high-pressure conditions. For instance, in high-performance chip cooling within micro-electromechanical systems, high-speed gas flowing through micro-nozzles generates localized jets on the chip surface. These jets enhance convective heat transfer by disrupting the thermal boundary layer through shear forces. Additionally, compressible effects at the nozzle exit generate expansion waves, reducing gas temperature, improving cooling efficiency, and mitigating shock wave effectsHo and Tai (1998); Stone, Stroock, and Ajdari (2004). In heap leaching with microporous seepage, high-pressure fluids moving through micropores and fractures induce local pressure and velocity fluctuations. These fluctuations enhance pore wall dissolution and facilitate fracture propagation. Shock waves at pore tips further assist in forming interconnected fracture pathwaysWang, Wang, and Chen (2018). These examples highlight the critical role of high-speed compressible flow in small-scale nonequilibrium processes, where the interplay of high pressures, localized effects, and compressibility drives significant physical and engineering effects.
@@
Due to the complexity of the Boltzmann equation’s collision term, developing simplified kinetic models that capture its most relevant features is an effective approach. Common simplified models include the Bhatnagar-Gross-Krook (BGK) modelBhatnagar, Gross, and Krook (1954), the ellipsoidal statistical BGK (ES-BGK) modelHolway (1966), the Shakhov modelShakhov (1968), and the Liu modelLiu (1990). To further enhance computational efficiency and applicability, researchers have explored various strategies for simplifying collision terms and tailoring models to specific problems. This has resulted in the development of various computational methods for solving cross-scale flow problems. Notable examples include the gas-kinetic schemeXu and Prendergast (1994); Xu (2001), unified gas-kinetic schemeXu and Huang (2010); Huang, Xu, and Yu (2012), discrete unified gas-kinetic schemeGuo, Xu, and Wang (2013); Xin, Zhang, and Guo (2023), discrete velocity methodYang et al. (2016); Yuan and Zhong (2020); Yang et al. (2022), and the lattice Boltzmann method (LBM)He and Luo (1997); Gan et al. (2008, 2011a); Xu et al. (2012); Wang, Wei, and Qian (2022); Shirsat, Nayak, and Patil (2022); Fei et al. (2023); Liang et al. (2024); Wang et al. (2017); Shan et al. (2024); Zhuang et al. (2024); Tao, Guo, and Wang (2017); Chen et al. (2024a); Wang, Wei, and Qian (2020); Wei et al. (2022); Fei, Derome, and Carmeliet (2024); Hou et al. (2025).
@@
Historically, the DBM evolved from the physical modeling branch of the LBM through selective taking, abandoning and supplementation. They start from the same point (kinetic theory), but aims at different requirements of the simulation studyXu, Zhang, and Gan (2024). As a discretized form of the simplified Boltzmann equation, DBM retains the use of discrete velocities but avoids relying on specific discrete formats, imposing only the essential physical constraints during discretization. These constraints include conserved kinetic moments and higher-order nonconserved kinetic moments, which are invariants in coarse-grained physical modeling and must be preserved. Such constraints provide a unique perspective for analyzing TNE behaviors, effects, and their manifestations. As the order and variety of higher-order kinetic moment constraints increase, DBM goes beyond the continuity assumption and near-equilibrium approximations of traditional fluid modeling. Although the model’s complexity increases slightly, its ability to describe nonequilibrium phenomena is significantly enhanced. Over time, DBM has developed new schemes for detecting, presenting, describing, and analyzing nonequilibrium states and effects based on the phase space description method and complex physical field analysis techniquesXu, Zhang, and Gan (2024). Specifically, each nonconserved kinetic moment provides a perspective on TNE intensity. These perspectives are interrelated and complementary, allowing for a more refined characterization of the physical structure and evolution of mesoscopic interfaces.
@@
While these studies advanced the exploration of TNEs in 3D cases, they are limited to situations with weak nonequilibrium effects. In scenarios involving large gradients or intense TNEs, these models often fail to capture the finer details of the nonequilibrium phenomena. To address this limitation, extending DBM to the 3D Burnett level is essential for achieving a more comprehensive characterization of nonequilibrium properties. In the field of higher-order DBM, Gan et al. derived expressions for nonlinear viscous stress and heat flux by considering second-order TNEs using Chapman-Enskog (CE) analysis. They constructed a Burnett-level DBM to investigate HNEs and TNEs in high-speed compressible flowsGan et al. (2018a). Later, they developed a multiscale, multiphase DBM based on density functional kinetic theory. This model, corresponding to the super-Burnett level, is suitable for studying TNEs in systems ranging from continuum to transition flow regimesGan et al. (2022).
Building upon previous work, this study constructs a 3D Burnett-level DBM to investigate HNE and TNEs in high-speed compressible flows. The core concept involves applying CE multiscale analysis to identify the kinetic moment relationships that should be preserved in coarse-grained physical modeling when considering second-order TNEs. Additionally, this study provides a theoretical derivation of nonlinear constitutive relations for the 3D case, analyzing how spatial dimensions influence the types of nonequilibrium driving forces and their coupling effects. To facilitate this, a suitable discrete velocity set is designed to discretize the phase space, enabling the development of a robust and efficient 3D DBM. The structure of the paper is organized as follows: Section II outlines the construction of the multiscale model. Section III validates the model’s accuracy in capturing large-scale structures in high-speed compressible flows, using standard test cases from 1D to 3D scenarios. Section IV evaluates the model’s ability to describe higher-order TNEs across scales. Finally, Section V concludes with a summary of the key findings.
@@ II Three-dimensional high-order discrete Boltzmann model
(2) Analysis stage: By utilizing higher-order nonconserved kinetic moments $(f-f{{}^{(0)}})$, DBM enables the efficient extraction of nonequilibrium states, effects, and behaviors. Moreover, constructing the phase space through the independent component of $\mathbf{M}_{m,n}(f-f{{}^{(0)}})$ offers a more intuitive representation of nonequilibrium states, effects, and behaviors.
@@ II.1 Simplification of the Boltzmann equation
In kinetic theory, the properties of a system are described by the distribution function and its kinetic moments. After discretization, the distribution function itself loses physical significance, while the integral or summation forms of the corresponding kinetic moments retain their physical relevance. Therefore, whether simplifying the collision term or discretizing the particle velocity space, the DBM must ensure that the fundamental physical constraint describing the system’s behavior, represented by the kinetic moments, remains invariant, both before discretization in the integral form and after discretization in the summation form, as expressed below:
@@
The CE multiscale analysis demonstrates that, recovering the hydrodynamic equations at different levels, requires satisfying specific kinetic moments. The DBM operates independently of the CE expansion. In DBM modeling, the CE multiscale analysis primarily aids in identifying and validating the kinetic moment relations retained in the coarse-grained model. The explicit derivation of macroscopic hydrodynamic equations is not a prerequisite for DBM modeling and simulations. To determine the required kinetic moments, the CE expansion is applied to both sides of Eq. (4), beginning with a multiscale expansion of the distribution function, time derivatives, and spatial derivatives:
@@
where $f_{i}^{(j)}$ represents the $j$th order deviation from $f_{i}^{(0)}$, and $\partial_{t_{j}}$ denotes the $j$th order term in the time scale expansion.
@@
It is evident that both $f_{i}^{(1)}$ and $f_{i}^{(2)}$ can be expressed in terms of $f_{i}^{(0)}$. Specifically, $f_{i}^{(1)}$ is represented by a first-order polynomial in $\mathbf{v}_{i}$, while $f_{i}^{(2)}$ is represented by a second-order polynomial in $\mathbf{v}_{i}$. Consequently, to satisfy the kinetic moment relations for $f_{i}^{(0)}$, the required accuracy must increase by one order for $f_{i}^{(1)}$ and by two orders for $f_{i}^{(2)}$. Thus, Eq. (5) can be rewritten as:
@@
$\mathbf{M}_{2,0}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})=\frac{1}{2}\rho[(n+3)RT+u^{2}],$ (14)
@@
$\mathbf{M}_{3,1}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}=\frac{1}{2}\rho\mathbf{u}[(n+5)RT+u^{2}],$ (16)
@@
$\mathbf{M}_{4,2}=$ $\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}\mathbf{v}_{i}=\frac{1}{2}\rho[(n+5)R^{2}T^{2}$ (18)
@@
$\mathbf{M}_{5,3}=\sum_{i}\frac{1}{2}f_{i}^{(0)}(v_{i}^{2}+\eta_{i}^{2})\mathbf{v}_{i}\mathbf{v}_{i}\mathbf{v}_{i}=\rho\left[\left(\frac{n+9}{2}RT+\frac{u^{2}}{2}\right)\mathbf{uuu}\right.$ (20)
@@
$\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{2,0},\mathbf{M}_{2},\mathbf{M}_{3,1},\mathbf{M}_{3},\mathbf{M}_{4,2},\mathbf{M}_{4},\mathbf{M}_{5,3})^{T}=(M_{0},M_{1x},M_{1y},M_{1z},\cdots,M_{5,3yzz},M_{5,3zzz})^{T}$ is a $55\times 1$ column matrix. Each element of $\bm{\Phi}$ represents an independent kinetic moment of $f_{i}^{(0)}$. The matrix $\mathbf{C}=(\mathbf{c}_{1},\mathbf{c}_{2},\cdots,\mathbf{c}_{55})$ is a $55\times 5$ 55 square matrix that establishes the relationship between the discrete equilibrium distribution function and the kinetic moments. Here, $\mathbf{c}_{i}$ is the $i$th column of $\mathbf{C}$, defined as: $\mathbf{c}_{i}=[1,v_{ix},v_{iy},v_{iz},\cdots,\frac{1}{2}(v_{i}^{2}+\eta_{i}^{2})v_{iz}v_{iz}v_{iz}]^{T}$, and $\mathbf{f}^{(0)}=(f_{1}^{(0)},f_{2}^{(0)},f_{3}^{(0)},\cdots,f_{55}^{(0)})^{T}$ represents the vector of discrete equilibrium distribution functions. Thus, $f_{i}^{(0)}$ can be computed as follows Gan et al. (2013):
@@
To efficiently handle the continuity of velocity space in numerical simulations, the originally continuous velocity space is discretized into a finite set of discrete velocities. The design principles and optimization guidelines of the discrete velocity sets (DVSs) can be elaborated from the following four aspectsGan et al. (2018b, 2022); He et al. (2025): (1) Determination of the number of discrete velocities: The minimum required number of discrete velocities is determined by the independent kinetic moment constraints that the distribution function must satisfy. In this study, these nine kinetic moments lead to 55 component equations, requiring at least 55 discrete velocities for accurate modeling. Increasing the number of discrete velocities improves the accuracy and stability of the model but reduces computational efficiency.
(2) Symmetry and asymmetry optimization: The discrete velocity model is constructed based on symmetry principles, including geometric and physical symmetry. Geometric symmetry ensures that discrete velocities are symmetrically distributed in space. Physical symmetry refers to the invariance of kinetic moment constraints under spatial rotation, requiring the moment matrix $\mathbf{C}$, constructed from discrete velocities, to be invertible. Consequently, a few asymmetric discrete velocities (e.g., velocities 52-55 in this study) are introduced.
(3) Discretization optimization strategies: The effectiveness of DVSs can be evaluated by comparing numerical and analytical solutions for nonequilibrium effects. A well-designed DVS produces smooth and continuous $f_{i}$ while efficiently utilizing its more components for maintaining numerical robustness.He et al. (2025)
(4) Flexibility in discrete velocity model design: DBM does not impose a fixed discrete velocity format. In practice, the DVS should be adjusted based on the problem characteristics and numerical feedback for optimal performance.
To ensure computational stability and accuracy, a 3D discrete velocity set with fifty-five velocities (D3V55), as shown in Fig. 1(a), is proposed. The mathematical expression is as follows:
$(v_{ix},v_{iy},v_{iz})=\begin{cases}(0,0,0)&\text{for }i=1,\\ cyc:c(\pm 1,0,0)&\text{for }2\leq i\leq 7,\\ cyc:c(\pm 1,\pm 1,0)&\text{for }8\leq i\leq 19,\\ c(\pm 1,\pm 1,\pm 1)&\text{for }20\leq i\leq 27,\\ cyc:c(\pm 1,\pm 2,0)&\text{for }28\leq i\leq 51,\\ c(3,-2,-1)\text{ or }c(-2,1,2)&\text{for }i=52,\\ c(-3,2,1)\text{ or }c(-2,3,-2)&\text{for }i=53,\\ c(2,1,-3)\text{ or }c(1,-2,-1)&\text{for }i=54,\\ c(-2,-1,3)\text{ or }c(3,-2,1)&\text{for }i=55.\end{cases}$ (23)
Here, "$cyc$" represents cyclic permutation, which generates the corresponding vector by cyclically rearranging the digits and signs in the velocity vector. The decomposition diagram is shown in Fig. 1 (b)-(f). Specifically, (i) the first velocity corresponds to the zero velocity; (ii) the 2nd to 7th velocities correspond to the face-centered velocities; (iii) the eighth to nineteenth velocities correspond to the second diagonal velocities; (iv) the twentieth to twenty-seventh velocities correspond to the cube vertex velocities; (v) the twenty-eighth to fifty-first velocities correspond to the third diagonal velocities; and (vi) the final four velocities are special velocities with insufficient symmetry, used to ensure the invertibility of the velocity matrix. We also set
$\eta_{i}=\begin{cases}10\eta_{0}&\text{for }i=1,\\ \eta_{0}&\text{for }2\leq i\leq 7,\,i=15,\,19,\,26,\\ &\hskip 13.00005pt35\leq i\leq 40,\\ \eta_{0}\text{ or }\eta_{i}=0&\text{for }52\leq i\leq 53,\\ 0&\text{for all other }i.\end{cases}$ (24)